OpenStudy (anonymous):

Simplify the expression below. If your answer is not a whole number, enter it as a fraction in lowest terms, using the slash mark ( / ) as the fraction bar. √7/4 x √9/7

OpenStudy (anonymous):

\[\frac{ \sqrt{7} }{ 4 } * \frac{ \sqrt{9} }{ 7 } = \frac{ \sqrt{7} }{ 4 } * \frac{ 3 }{ 4 } = \frac{ 3\sqrt{7} }{ 16 }\]

OpenStudy (anonymous):

Salut Mohamed ! ;) Your expression is : \[\sqrt\frac74\times \sqrt\frac97\] or what ?

OpenStudy (anonymous):

Wow, if the expression is what Noura11 said what i did it's wrong.

OpenStudy (anonymous):

yeah it was 3/2

OpenStudy (anonymous):

solve for x log 16 x =3/2

OpenStudy (anonymous):

\[\sqrt{\frac{ 7 }{ 4 }} * \sqrt{\frac{ 9 }{ 7 }} = \sqrt{\frac{ 7*9 }{ 4*7 }} = \sqrt{\frac{ 9 }{ 4 }} = \frac{ \sqrt{9} }{ \sqrt{4} } = \frac{3}{2}\]

OpenStudy (anonymous):

wha????

OpenStudy (anonymous):

thanks that's the answer but can u solve the upper one is a logarithm solve for x

OpenStudy (anonymous):

We have : \[ \log_{16} x =3/2\iff x=16^{\frac32} \\\iff x=\sqrt{16^3} =16\sqrt{16}=16\times4=64\]